8 A $20,000 face value STRIPS is currently quoted at 38.642 and has 8 years to...

8

A $20,000 face value STRIPS is currently quoted at 38.642 and has 8 years to maturity. What is the yield-to-maturity?

12.83 percent

6.26 percent

6.30 percent

12.65 percent

12.25 percent
Answer #1

Answer is 12.25%

Face Value = $20,000

Current Price = 38.642% * Face Value
Current Price = 38.642% * $20,000
Current Price = $7,728.40

Time to Maturity = 8 years
Semiannual Period = 16

Let Semiannual YTM be i%

$7,728.40 = $20,000 * PVIF(i%, 16)
$7,728.40 = $20,000 / (1 + i)^16
(1 + i)^16 = 2.5878578
1 + i = 1.06123
i = 0.06123 or 6.123%

Semiannual YTM = 6.123%
Annual YTM = 2 * 6.123%
Annual YTM = 12.246% or 12.25%